OpenStudy (anonymous):

test for convergence integrate (0 to pi/2) cotx dx

OpenStudy (raden):

integral by sub.. int (cot(x)) dx = int (cos(x)/sin(x)) dx let u=sin(x) du=cos(x) dx, so the integration can be int (1/u) du = ln(u) subtitute back that u=sin(x) = ln(sin(x)) now, just calculate for limits integral from 0 to pi/2